# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 4.1.1 The model
$Z=\int[d{\bm{\Phi}}]\,[d{\bm{A}}]\,e^{-S({\bm{\Phi}},{\bm{A}})},\quad S({\bm{\Phi}},{\bm{A}})=\int d^{d}x\,\left[{\cal L}_{\rm AH}({\bm{\Phi}},{\bm{A}})+{\cal L}_{\rm gf}({\bm{A}})\right],\quad{\cal L}_{\rm gf}({\bm{A}})={1\over 2\lambda}(\partial_{\mu}A_{\mu})^{2},$ (22)
@@
The RG flow of the AHFT has been studied using the perturbative $\varepsilon\equiv 4-d$ expansion [18, 273, 274, 54, 83] and the functional RG approach  [81]. The analysis of the corresponding fixed points (FPs) allows one to determine the nature of the transitions in this class of systems. In particular, if a stable FP exists, continuous transitions are possible. The model has also been studied using the large-$N$ expansion, obtaining the expansion of some critical exponents for large $N$ [18, 205, 55, 64, 76].
@@
where $y_{q}$ is RG dimension of the operator $Q$, and $\eta_{q}$ is the usual critical exponent. The approach to the critical point is controlled by the length-scale critical exponent $\nu$, which can be defined from the power-law divergence of the length scale $\xi$ of the critical modes, i.e. $\xi\sim|r-r_{c}|^{-\nu}$, where $r$ is the coefficient of the quadratic term in Eq. (18) and $r_{c}$ its critical value.
@@ 4.1.2 Renormalization-group flow in the perturbative $\varepsilon$-expansion
The RG flow of the SU($N$)-symmetric AHFT field theory has been determined close to four dimensions within the $\varepsilon$-expansion framework. The RG scaling functions that are related with the diverging renormalization constants have been determined perturbatively (perturbative calculations have been performed in the Lorenz gauge), in powers of $\varepsilon\equiv 4-d$ [94, 18]. In the perturbative calculations dimensional regularization and the minimal-subtraction (MS) renormalization scheme are generally used, see, e.g., Refs. [2, 95]. The corresponding MS $\beta$ functions associated with the renormalized couplings $\alpha\equiv g^{2}$ and $u$ have been computed to four loops [83]. The FPs of the RG flow are obtained from the common zeroes of the $\beta$ function. Their stability is controlled by the eigenvalues of the matrix $\Omega_{ij}=\partial_{g_{i}}\beta_{j}$ (where $g_{i}$ indicates the Lagrangian couplings, and $\beta_{i}$ the corresponding $\beta$ function) [2, 95]: a FP is stable if all eigenvalues $\lambda_{i}$ of $\Omega$, computed at the FP, are positive.
@@
We expect that the value $N^{*}$ above which the theory has a stable FP depends on the dimension $d$, i.e., that, in generic dimension $d$, a stable FP exists only for $N>N^{*}(d)$, with $N^{*}(4)=N_{4}$. Therefore, 3D AH models may undergo a charged continuous transition only if $N>N^{*}(3)$. The critical number of components $N^{\star}(d)$ has been determined to four loops [83]:
@@
The determination of $N^{*}(3)$ from the asymptotic expansion (30) is quite difficult because of the large coefficients. Moreover, the perturbative expansions for gauge theories are not Borel summable, because of the presence of renormalons, see, e.g., Refs. [275, 276, 277, 278, 279], so that it is not clear whether the resummation methods  [2] that are successfully used for $\Phi^{4}$ theories [^18] The $\varepsilon$-expansion for $\Phi^{4}$ theories is conjectured to be Borel summable [2]. may provide reasonably accurate estimates. In spite of these difficulties, by means of a resummation method that makes the reasonable assumption that $N^{*}(d)$ vanishes linearly for $d\to 2$ (this is suggested by the $D=2+\varepsilon$ perturbative results of Refs. [280, 281]), Ref. [83] obtained $N^{*}(3)=12(4)$ in three dimensions, which confirms the absence of a stable FP—and therefore, of continuous transitions—for small values of $N$. The field-theoretical estimate is in reasonable agreement with the numerical estimate [87, 90] $N^{*}(3)=7(2)$ obtained in MC simulations of LAH models, see also Sec. 5.3.  [^19] A similar behavior has been observed  [200] in the quantum square-lattice SU($N$) antiferromagnets at the transition between the SU($N$) Néel phase and the valence-bond solid phase. The numerical FSS analyses of the Rényi entanglement entropy reported in Ref. [200] are apparently compatible with the CFT predictions appropriate for continuous transitions only for $N\geq 8$, leading to the conclusions that systems with $N\leq 7$ undergo (weak) first-order transitions. The $N$-component AHFT (22) is a candidate field theory for the critical behavior at these transitions, see, e.g., Refs. [27, 29, 28, 30, 200]. If this is the case, the results of Ref. [200] also confirm the absence of a FP for small values of $N$.
It is interesting to observe that, to all order of perturbation theory, the $\beta$ function of the gauge coupling $\alpha$ can be expressed in terms of the anomalous dimension $\eta_{A}$ of the scalar field in the Lorenz gauge [53, 2]:
$\beta_{\alpha}=\alpha(d-4+\eta_{A}).$ (31)
Therefore, the fixed-point condition $\beta_{\alpha}=0$ implies at the charged FP ($\alpha^{*}\not=0$) the exact relation [53, 19]
$\eta_{A}=4-d.$ (32)
@@ 4.1.3 Three-dimensional Abelian Higgs field theory in the large-$N$ limit
In the presence of a gauge fixing we can also consider correlation functions that are not gauge invariant. For example, we may consider the correlation function $G_{\Phi}({\bm{x}},{\bm{y}})\equiv\langle\bar{\bm{\Phi}}({\bm{x}})\cdot{\bm{\Phi}}({\bm{y}})\rangle$ of the Lagrangian field ${\bm{\Phi}}({\bm{x}})$. In the Lorenz gauge the anomalous dimension of ${\bm{\Phi}}({\bm{x}})$ is given by [18, 76]
$\eta_{\phi}=-{20\over\pi^{2}}{1\over N}+O(N^{-2}),$ (36)
which allows one to compute the RG dimension of $\Phi$ in the Lorenz gauge, $y_{\phi}=(1+\eta_{\phi})/2$.
@@ 4.2.1 SO($N$)-symmetric Abelian Higgs field theories: Definition of the model
To determine the possible ordered phases of the model with $N\geq 2$, one can use the mean-field approximation, since field fluctuations are only relevant along the transition lines. This amounts to analyzing the minima of the scalar potential  [282]. One finds two ordered phases, with different symmetry, depending on the sign of $v$. For $v<0$, the ordered phase is invariant under SO($N-1$) transformations, so that the symmetry-breaking pattern at the transition is the same as that of the O($N$) vector model,
@@
4.2.2 SO($N$)-symmetric Abelian-Higgs field theories: Renormalization-group flow
To determine the possible charged transitions in SO($N$) invariant models, one can study the RG flow in the MS renormalization scheme close to four dimensions, as was done for the SU($N$) symmetric AHFT in Sec. 4.1.2. The one-loop $\beta$ functions associated with the renormalized couplings $\alpha$, $u$, and $v$, are given by [273, 274]
@@ 4.2.2 SO($N$)-symmetric Abelian-Higgs field theories: Renormalization-group flow
Note that $v^{*}>0$, so that this FP is only relevant for charged transitions characterized by the symmetry-breaking pattern given in Eq. (39). There are no higher-order computations, that allow one to determine $N_{o}^{*}(d)$ in $d$ dimensions and, in particular, for $d=3$. However, by analogy with the SU($N$) case, one expects $N_{o}^{*}(d=3)$ to be of order ten  [282].
@@ 4.2.3 Further extensions of the Abelian Higgs field theories
${\cal L}_{P}=\frac{1}{4g^{2}}\,F_{\mu\nu}^{2}+\overline{D_{\mu}{\bm{\Phi}}}\cdot D_{\mu}{\bm{\Phi}}+r\,\bar{\bm{\Phi}}{\bm{\Phi}}+u\,(\bar{\bm{\Phi}}\cdot{\bm{\Phi}})^{2}+v\,|{\bm{\Phi}}\cdot{\bm{\Phi}}|^{2}+w\,\sum_{a=1}^{N}(\bar{\Phi}^{a}\Phi^{a})^{2},$ (44)
@@ 5.1.1 Definition of the model
${\bm{z}}_{\bm{x}}\to e^{i\Lambda_{\bm{x}}}{\bm{z}}_{\bm{x}},\qquad A_{{\bm{x}},\mu}\to A_{{\bm{x}},\mu}+\Lambda_{\bm{x}}-\Lambda_{{\bm{x}}+\hat{\mu}},$ (47)
with $\Lambda_{\bm{x}}\in\mathbb{R}$. Model (45) can be seen as a particular, quite straightforward, lattice regularization of the AHFT defined by the Lagrangian (18).
@@
We will also consider the Lorenz gauge fixing, defined by requiring [286]
@@
for all lattice sites $\bm{x}$. In the presence of C boundary conditions, the above Lorenz gauge condition provides a complete gauge fixing as it completely breaks gauge invariance and thus it makes gauge correlations well defined. As demonstrated in Ref. [286], the lattice Lorenz gauge fixing (49) has an important property. When computed in the Lorenz gauge, Fourier-transformed gauge correlations are also well defined in the infinite-volume limit for any nonvanishing momentum, at variance with what happens in other gauges. For instance, axial gauges suffer from the existence of an infinite two-dimensional family of quasizero modes for $L$ large. These zero modes give rise to spurious divergences, unrelated to the presence of long-range physical correlations [286]. Note that, in perturbation theory, the gauge-fixed theory is obtained by adding a term $\alpha\sum_{\mu}(\Delta_{\mu}^{-}A_{{\bm{x}},\mu})^{2}$ to the Hamiltonian. As discussed in Ref. [286] this soft gauge fixing is not appropriate for the nonperturbative analysis of the phase behavior of the LAH model, as it suffers from the presence of propagating unphysical longitudinal modes, which may hide the physical signal. Similar conclusions are reached from the analysis of scalar correlation functions, see Ref. [91].
@@
The unitary gauge fixing is not complete, indeed the Hamiltonian $H_{\rm ug}$ is still invariant under gauge transformations in which $\Lambda_{\bm{x}}$ is a multiple of $2\pi$. As discussed in Sec. 3.5.3, this residual gauge symmetry can induce topological transitions: The model (50) represents a soft version of the inverted XY (IXY) gauge model defined in Sec. 3.5.1, which is obtained for $J\to\infty$. Indeed, in this limit, we have $\cos A_{{\bm{x}},\mu}=1$, so that $A_{{\bm{x}},\mu}$ takes only values that are multiples of $2\pi$ modulo gauge transformations, i.e., $A_{{\bm{x}},\mu}=2\pi n_{{\bm{x}},\mu}$ with $n_{{\bm{x}},\mu}\in\mathbb{Z}$.
The qualitative features of the phase diagram of the model depend on $N$, as sketched in Fig. 2. For $N=1$ there are only two phases, which differ in the topological properties of the gauge fields and in the confinement properties of the charged nonlocal excitations. For $N>1$ the phase diagram is more complex, due to the possibility of the spontaneous breaking of the global SU($N$) symmetry. Note that a Higgs phase is always present for large $J$ and $\kappa$.
Also for $N>1$ a line of topological transitions is present, and the mechanism underlying these transitions is the one discussed in Sec. 3.5.3: in the broken-symmetry phase the minima of the potential correspond (in the infinite volume limit and up to gauge transformations) to ${\bm{z}}_{\bm{x}}={\bm{z}}_{{\bm{x}}+\hat{\mu}}$ and $\lambda_{{\bm{x}},\mu}=1$, and these minima are still invariant under gauge transformations in which $\Lambda_{\bm{x}}$ is a multiple of $2\pi$, exactly as in $N=1$ case.
Figure 2: The $\kappa$-$J$ phase diagram of the $N$-component LAH model (45), for $N=1$ (left), and generic $N\geq 2$ (right). For $N=1$, there are two phases, the Coulomb (C) and Higgs (H) phases, characterized by the confinement and deconfinement of charged gauge-invariant excitations, respectively. For $N\geq 2$, the scalar field is disordered and gauge correlations are long ranged in the small-$J$ Coulomb (C) phase. For large $J$ two phases occur, the molecular (M) and Higgs (H) ordered phases, in which the global SU($N$) symmetry is spontaneously broken. The two phases are distinguished by the behavior of the gauge modes: the gauge field is long ranged in the M phase (small $\kappa$), while it is gapped in the H phase (large $\kappa$). Moreover, while the C and M phases are confined phases, in the H phase charged gauge-invariant excitations are deconfined.

### v3

@@ 4.1.1 The model
$Z=\int[d{\bm{\Phi}}]\,[d{\bm{A}}]\,e^{-S({\bm{\Phi}},{\bm{A}})},\quad S({\bm{\Phi}},{\bm{A}})=\int d^{d}x\,\left[{\cal L}_{\rm AH}({\bm{\Phi}},{\bm{A}})+{\cal L}_{\rm gf}({\bm{A}})\right],\quad{\cal L}_{\rm gf}({\bm{A}})={1\over 2\zeta}(\partial_{\mu}A_{\mu})^{2},$ (22)
@@
The RG flow of the AHFT has been studied using the perturbative $\varepsilon\equiv 4-d$ expansion [18, 304, 305, 54, 83] and the functional RG approach  [81]. The analysis of the corresponding FPs allows one to determine the nature of the transitions in this class of systems. In particular, if a stable FP exists, continuous transitions are possible in systems with analogous features. The model has also been studied using the large-$N$ expansion, obtaining the expansion of some critical exponents for large $N$ [18, 234, 55, 64, 76].
@@
where $y_{q}$ is the RG dimension of the operator $Q$ and $\eta_{q}$ is the usual critical exponent. The approach to the critical point is controlled by the length-scale critical exponent $\nu$, which can be defined from the power-law divergence of the length scale $\xi$ of the critical modes, i.e. $\xi\sim|r-r_{c}|^{-\nu}$, where $r$ is the coefficient of the quadratic term in Eq. (18) and $r_{c}$ its critical value.
@@ 4.1.2 Renormalization-group flow in the perturbative $\varepsilon$-expansion
The RG flow of the SU($N$)-symmetric AHFT field theory has been determined close to four dimensions in the $\varepsilon$-expansion framework. The RG scaling functions that are related with the diverging renormalization constants have been determined perturbatively [perturbative calculations are usually performed using the Lorenz gauge fixing as in Eq. (22)], in powers of $\varepsilon\equiv 4-d$ [95, 18]. In the perturbative calculations dimensional regularization and the minimal-subtraction (MS) renormalization scheme are generally used, see, e.g., Refs. [2, 96]. The corresponding MS $\beta$ functions associated with the renormalized couplings $\alpha\equiv g^{2}$ and $u$ have been computed to four loops [83]. The FPs of the RG flow are obtained from the common zeroes of the $\beta$ function. Their stability is controlled by the eigenvalues of the matrix $\Omega_{ij}=\partial_{g_{i}}\beta_{j}$ (where $g_{i}$ indicates the Lagrangian couplings, and $\beta_{i}$ the corresponding $\beta$ function) [2, 96]: A FP is stable if all eigenvalues $\lambda_{i}$ of $\Omega$, computed at the FP, are positive.
@@
The value $N^{*}$ above which the theory has a stable FP depends on the dimension $d$, thus, in the $\varepsilon$ expansion approach one defines a function $N^{*}(d)$ such that a stable FP exists only for $N>N^{*}(d)$, with $N^{*}(4)=N_{4}$. 3D AH models may undergo a charged continuous transition only if $N>N^{*}(3)$. The critical number of components $N^{\star}(d)$ has been determined to four loops [83]:
@@
The determination of $N^{*}(3)$ from the asymptotic expansion (30) is quite difficult because of the large coefficients. Moreover, the perturbative expansions for gauge theories are not Borel summable, because of the presence of renormalons, see, e.g., Refs. [306, 307, 308, 309, 310], so it is not clear whether the resummation methods  [2] that are successfully used for $\Phi^{4}$ theories [^19] The $\varepsilon$-expansion for $\Phi^{4}$ theories is conjectured to be Borel summable [2]. may provide reasonably accurate estimates. In spite of these difficulties, by means of a resummation method that makes the reasonable assumption that $N^{*}(d)$ vanishes linearly for $d\to 2$ (this is suggested by the $D=2+\varepsilon$ perturbative results of Refs. [311, 312]), Ref. [83] obtained $N^{*}(3)=12(4)$ in three dimensions, which confirms the absence of a stable FP—and therefore, of continuous transitions—for small values of $N$. The field-theoretical estimate is in reasonable agreement with the estimate [87, 90] $N^{*}(3)=7(2)$ obtained by numerical analyses of the LAH models, see also Sec. 5.3.202020A similar behavior has been observed  [229] in quantum square-lattice SU($N$) antiferromagnets at the transition between the SU($N$) Néel phase and the valence-bond solid phase. The numerical FSS analyses of the Rényi entanglement entropy reported in Ref. [229] are apparently compatible with the CFT predictions appropriate for continuous transitions only for $N\geq 8$, leading to the conclusions that systems with $N\leq 7$ undergo (weak) first-order transitions. The $N$-component AHFT (22) is a candidate field theory for the critical behavior at these transitions, see, e.g., Refs. [27, 29, 28, 30, 229]. If this is the case, then the results of Ref. [229] confirm the absence of a FP for small values of $N$.
It is interesting to observe that, to all orders of perturbation theory, the $\beta$ function of the gauge coupling $\alpha$ can be expressed in terms of the anomalous dimension $\eta_{A}$ of the scalar field in the Lorenz gauge [53, 2, 313]:
$\beta_{\alpha}=\alpha\,[d-4+\eta_{A}(\alpha,u)],\qquad\eta_{A}(\alpha,u)=\frac{\mathrm{d}\log Z_{A}}{\mathrm{d}\log\mu},$ (31)
where $Z_{A}$ is the MS renormalization constant associated with the gauge field. Therefore, at the charged FP ($\alpha^{*}\not=0$) the FP condition $\beta_{\alpha}=0$ implies the exact relation [53, 19]
$\eta_{A}\equiv\eta_{A}(\alpha^{*},u^{*})=4-d.$ (32)
@@
4.1.3 Role of the gauge fixing
As already mentioned above, the Lorenz gauge-fixing term ${\cal L}_{\rm gf}$ in Eq. (22) is necessary to obtain a well-defined functional path integral, and, in particular, a well-defined perturbative expansion, see, e.g., Ref. [2]. Appropriate Ward identities guarantee that the gauge-invariant observables do not depend on the parameter $\zeta$ appearing in the gauge-fixing term ${\cal L}_{\rm gf}$. In particular, the MS perturbative $\beta$ functions $\beta_{\alpha}$ and $\beta_{u}$, the corresponding FPs, and the critical exponents are independent of $\zeta$. However, as discussed in Sec. 5.1, and, in more detail, in Refs. [46, 47, 48, 49, 50, 313, 91], the gauge fixing plays an important role also in lattice AH models. In particular, different behaviors are observed for positive values of $\zeta$ (in this case we speak of a soft Lorenz gauge fixing) and in the limiting case $\zeta=0$ (hard Lorenz gauge fixing). For instance, only in the second case the nonlocal gauge-invariant charge operators that condense in the low-temperature Higgs phase admit a local field representation [313, 91]. To better understand the role of the gauge fixing in the lattice theory, it is useful to discuss the RG flow of the Lorenz gauge parameter in the corresponding AHFT.
As a consequence of the Ward identities that follow from gauge invariance, which imply that the gauge-fixing term ${\cal L}_{\rm gf}$ in Eq. (22) does not renormalize, see, e.g., Ref. [2], the $\beta$ function of the parameter $\zeta$ takes the simple form
$\beta_{\zeta}\equiv\mu{\partial\zeta\over\partial\mu}=-\zeta\,\eta_{A}(\alpha,u).$ (34)
Eq. (34) shows that the value $\zeta=0$ represents a FP of the RG flow in the theory with a Lorenz gauge fixing. Moreover, since $\eta_{A}>0$, this FP is unstable. If we start the RG flow from $\zeta>0$, then $\zeta$ increases towards infinity, and the fluctuations of the nongauge-invariant modes are expected to become unbounded in this limit: nongauge-invariant correlations are expected to become disordered on large scales as in the absence of a gauge fixing. Therefore, we expect to be able to observe a critical behavior in the nongauge invariant sector only for $\zeta=0$, i.e., in the presence of a hard Lorenz gauge fixing. This is confirmed by the numerical results discussed in Refs. [313, 91].
@@ 4.1.4 Three-dimensional Abelian Higgs field theory in the large-$N$ limit
In the presence of a gauge fixing one can also consider correlation functions that are not gauge invariant. For example, we may consider the correlation function $G_{\Phi}({\bm{x}},{\bm{y}})\equiv\langle\bar{\bm{\Phi}}({\bm{x}})\cdot{\bm{\Phi}}({\bm{y}})\rangle$ of the Lagrangian field ${\bm{\Phi}}({\bm{x}})$. The anomalous dimension of ${\bm{\Phi}}({\bm{x}})$ depends on the gauge-fixing parameter $\zeta$ and is given by [18, 76]
$\eta_{\phi}=-{20+8\zeta\over\pi^{2}}{1\over N}+O(N^{-2}).$ (37)
As already discussed in Sec. 4.1.3, only in the presence of a hard Lorenz gauge fixing ($\zeta=0$) nongauge invariant correlations can become critical. In particular, as we shall see in Sec. 5.1.2, only in this case the Higgs phase can be characterized by the condensation of the scalar field, reflecting the condensation of a nonlocal gauge-invariant charged operator. For $\zeta=0$ we obtain
$\eta_{\phi}=-{20\over\pi^{2}}{1\over N}+O(N^{-2})\quad{\rm for}\;\;\zeta=0,$ (38)
which allows one to compute the RG dimension of $\Phi$ in the hard Lorenz gauge: $y_{\phi}=(1+\eta_{\phi})/2$.
@@ 4.2.1 SO($N$)-symmetric Abelian Higgs field theories: Definition of the model
To determine the possible ordered phases of the model with $N\geq 2$, one can use the mean-field approximation, since field fluctuations are only relevant along the transition lines. This amounts to analyzing the minima of the scalar potential  [314]. One finds two ordered phases, with different symmetry, depending on the sign of $v$. For $v<0$, the ordered phase is invariant under SO($N-1$) transformations, so the symmetry-breaking pattern at the transition is the same as that of the O($N$) vector model,
@@
4.2.2 SO($N$)-symmetric Abelian Higgs field theories: Renormalization-group flow
To determine the possible charged transitions in SO($N$) invariant models, one can study the RG flow in the MS renormalization scheme close to four dimensions, as it was done for the SU($N$) symmetric AHFT in Sec. 4.1.2. The one-loop $\beta$ functions associated with the renormalized couplings $\alpha$, $u$, and $v$, are given by [304, 305]
@@ 4.2.2 SO($N$)-symmetric Abelian Higgs field theories: Renormalization-group flow
Note that $v^{*}>0$, so this FP is only relevant for charged transitions characterized by the symmetry-breaking pattern given in Eq. (41). There are no higher-order computations, that allow one to determine $N_{o}^{*}(d)$ in $d$ dimensions and, in particular, for $d=3$. However, by analogy with the SU($N$)-symmetric case, one expects $N_{o}^{*}(d=3)$ to be of order ten  [314].
@@ 4.2.3 Further extensions of the Abelian Higgs field theories
${\cal L}_{P}=\frac{1}{4g^{2}}\,F_{\mu\nu}^{2}+\overline{D_{\mu}{\bm{\Phi}}}\cdot D_{\mu}{\bm{\Phi}}+r\,\bar{\bm{\Phi}}\cdot{\bm{\Phi}}+u\,(\bar{\bm{\Phi}}\cdot{\bm{\Phi}})^{2}+v\,|{\bm{\Phi}}\cdot{\bm{\Phi}}|^{2}+w\,\sum_{a=1}^{N}(\bar{\Phi}^{a}\Phi^{a})^{2},$ (46)
@@ 5 Lattice Abelian Higgs models with noncompact gauge variables
We present their phase diagrams, with several phases, including Higgs phases, and their critical behavior along the transition lines separating the different phases. Different scenarios emerge in the case of one-component and multi-component scalar fields. The observed critical behaviors can be interpreted using the approaches outlined in Sec. 3. In particular, the one-component models undergo topological transitions belonging to the IXY universality class, see Sec. 3.5.1, separating their Coulomb and Higgs phases. Multi-component models show a more complex phase diagram, arising from the nontrivial interplay between the global SU($N$) or SO($N$) symmetry and the local U(1) gauge symmetry. In particular, for a sufficiently large number of scalar components the transitions along the line that separates the Coulomb and Higgs phases are continuous and associated with the stable FP of the RG flow of the corresponding AHFT. We also emphasize the role of the Lorenz gauge fixing, which allows one to probe the condensation of extended charged excitations in the Higgs phase.
@@ 5.1.1 Definition of the model
${\bm{z}}_{\bm{x}}\to e^{i\Lambda_{\bm{x}}}{\bm{z}}_{\bm{x}},\qquad A_{{\bm{x}},\mu}\to A_{{\bm{x}},\mu}+\Lambda_{\bm{x}}-\Lambda_{{\bm{x}}+\hat{\mu}},\qquad\Lambda_{\bm{x}}\in\mathbb{R}.$ (49)
Model (47) can be seen as a particular, quite straightforward, lattice regularization of the AHFT defined by the Lagrangian (18).
@@
As we shall see, some properties of the model can be best analyzed by adding a Lorenz gauge fixing. It is defined by requiring [313]
@@
for all lattice sites $\bm{x}$. In the presence of C boundary conditions, the above Lorenz gauge condition represents a complete gauge fixing as it completely breaks gauge invariance and thus it makes gauge correlations well defined. As demonstrated in Ref. [313], the lattice Lorenz gauge fixing (51) has an important property. When computed in the Lorenz gauge, Fourier-transformed gauge correlations are also well defined in the infinite-volume limit for any nonvanishing momentum, at variance with what happens in other gauges. For instance, axial gauges suffer from the existence of an infinite two-dimensional family of quasizero modes for $L$ large. These zero modes give rise to spurious divergences, unrelated to the presence of long-range physical correlations [313].
It is important to note that in the AHFT with action (22) gauge invariance is broken by adding ${\cal L}_{\rm gf}\sim\zeta^{-1}(\partial_{\mu}A_{\mu})^{2}$ to the gauge-invariant Lagrangian. In the lattice case, this corresponds to adding
${1\over 2\zeta}\sum_{{\bm{x}},\mu}(\Delta_{\mu}^{-}A_{{\bm{x}},\mu})^{2}$ (52)
to the lattice Hamiltonian, Eq. (47). We name this gauge-symmetry breaking a soft Lorenz gauge fixing. The gauge fixing condition (51) (hard Lorenz gauge fixing) is straightforwardly recovered in the $\zeta\to 0$ limit. As discussed in Ref. [313] soft gauge fixings are not appropriate for the nonperturbative analysis of the phase behavior of the LAH model, as they suffer from the presence of propagating unphysical longitudinal modes, which may hide the physical signal. Similar conclusions are reached from the analysis of scalar correlation functions, see Ref. [91], and from the RG flow of the Lorenz parameter $\zeta$ in field theory, see Sec. 4.1.3.
@@
The unitary gauge fixing is not complete, indeed the Hamiltonian $H_{\rm ug}$ is still invariant under gauge transformations in which $\Lambda_{\bm{x}}$ is a multiple of $2\pi$. As discussed in Sec. 3.5.3, this residual gauge symmetry can induce topological transitions: The model (53) represents a soft version of the inverted XY (IXY) gauge model defined in Sec. 3.5.1, which is obtained for $J\to\infty$. Indeed, in this limit, we have $\cos A_{{\bm{x}},\mu}=1$, so $A_{{\bm{x}},\mu}$ takes only values that are multiples of $2\pi$, i.e., $A_{{\bm{x}},\mu}=2\pi n_{{\bm{x}},\mu}$ with $n_{{\bm{x}},\mu}\in\mathbb{Z}$.
The qualitative features of the phase diagram of the model depend on $N$, as sketched in Fig. 2. For $N=1$ there are only two phases, which differ in the topological properties of the gauge fields and in the confinement properties of the charged nonlocal excitations. For $N\geq 2$ the phase diagram is more complex, due to the possibility of the spontaneous breaking of the global SU($N$) symmetry. Note that a Higgs phase is always present for large $J$ and $\kappa$.
For $N>1$ a line of topological transitions (it separates the M and H phases) is also present. The mechanism underlying these transitions is the one discussed in Sec. 3.5.3: in the broken-symmetry phase the minimum-energy configurations correspond (in the infinite-volume limit and up to gauge transformations) to ${\bm{z}}_{\bm{x}}={\bm{z}}_{{\bm{x}}+\hat{\mu}}$ and $\lambda_{{\bm{x}},\mu}=1$. They are still invariant under gauge transformations in which $\Lambda_{\bm{x}}$ is a multiple of $2\pi$, exactly as for $N=1$, so we expect topological transitions analogous to the CH transitions for $N=1$.
Figure 2: The $\kappa$-$J$ phase diagram of the $N$-component LAH model (47), for $N=1$ (left), and generic $N\geq 2$ (right). For $N=1$, there are two phases, the Coulomb (C) and the Higgs (H) phase, characterized by the confinement and deconfinement of charged gauge-invariant excitations, respectively. For $N\geq 2$, the scalar field is disordered and gauge correlations are long ranged in the small-$J$ Coulomb (C) phase. For large $J$ two phases occur, the molecular (M) and Higgs (H) ordered phases, in which the global SU($N$) symmetry is spontaneously broken. The two phases are distinguished by the behavior of the gauge modes: the gauge field is long ranged in the M phase (small $\kappa$), while it is gapped in the H phase (large $\kappa$). Moreover, while the C and M phases are confined phases, in the H phase charged gauge-invariant excitations are deconfined.